# SNN4Agents: A Framework for Developing Energy-Efficient Embodied Spiking Neural Networks for Autonomous Agents

### v1

@@ Abstract
Recent trends have shown that autonomous agents, such as Autonomous Ground Vehicles (AGVs), Unmanned Aerial Vehicles (UAVs), and mobile robots, effectively improve human productivity in solving diverse tasks. However, since these agents are typically powered by portable batteries, they require extremely low power/energy consumption to operate in a long lifespan. To solve this challenge, neuromorphic computing has emerged as a promising solution, where bio-inspired Spiking Neural Networks (SNNs) use spikes from event-based cameras or data conversion pre-processing to perform sparse computations efficiently. However, the studies of SNN deployments for autonomous agents are still at an early stage. Hence, the optimization stages for enabling efficient embodied SNN deployments for autonomous agents have not been defined systematically. Toward this, we propose a novel framework called SNN4Agents that consists of a set of optimization techniques for designing energy-efficient embodied SNNs targeting autonomous agent applications. Our SNN4Agents employs weight quantization, timestep reduction, and attention window reduction to jointly improve the energy efficiency, reduce the memory footprint, optimize the processing latency, while maintaining high accuracy. In the evaluation, we investigate use cases of event-based car recognition, and explore the trade-offs among accuracy, latency, memory, and energy consumption. The experimental results show that our proposed framework can maintain high accuracy (i.e., 84.12% accuracy) with 68.75% memory saving, 3.58x speed-up, and 4.03x energy efficiency improvement as compared to the state-of-the-art work for NCARS dataset, thereby enabling energy-efficient embodied SNN deployments for autonomous agents.
@@ 2 Introduction
In recent years, the interest in implementing neuromorphic artificial intelligence based on Spiking Neural Networks (SNNs) for autonomous agents (so-called SNN-based autonomous agents) has rapidly increased. The reason is that, SNNs can offer high accuracy due to effective learning mechanism (Rathi et al., 2023; Putra and Shafique, 2021b), low computation latency due to efficient neural/spike coding (Guo et al., 2021), and ultra low power/energy consumption due to sparse spike-based operations (Putra and Shafique, 2020). To realize such systems in real life, capabilities of solving machine learning (ML) tasks like image classification (Putra et al., 2021, 2022a, 2023), object detection (Viale et al., 2021; Cordone et al., 2022), or object segmentation (Li et al., 2022) from images/videos are required. Besides such functionalities, SNN-based autonomous agents also require (1) small memory footprint as they typically employ resource-constrained hardware platforms, (2) low power/energy consumption to preserve the battery lifespan as they are typically powered by portable batteries, and (3) real-time output with high accuracy to provide quick decision (Bonnevie et al., 2021; Putra and Shafique, 2022, 2023b). To maximize the benefits of SNN sparse operations, event-based data can be employed as it directly provides a compatible data format for SNN processing and minimizes the pre-processing stage, such as the data-to-spike conversion (e.g., pixel data to spike train) and the spike coding. Therefore, the developments of SNN-based autonomous agents also need to consider event-based data, such as the NCARS dataset (Sironi et al., 2018; Putra et al., 2024; Bano et al., 2024).
@@ 2.1 State-of-the-Art Works and Their Limitations
The study for developing SNN-based autonomous agents is still at an early stage. The state-of-the-art works are summarized in Table 1. Here, we observe that most of the existing works focus on proposing frameworks and/or techniques for achieving high accuracy (Putra and Shafique, 2022, 2023a, 2023b, 2024). However, these works have not considered event-based data workload, therefore requiring a relatively complex pre-processing stage, including data-to-spike conversion and spike coding. Some other works explore techniques for achieving high accuracy considering event-based automotive data (Viale et al., 2021, 2022; Cordone et al., 2022). However, these works have not considered optimizing the model size to fit into resource-constrained autonomous agents. All the above-discussed limitations of the state-of-the-art expose that, the optimization stages for enabling efficient SNN deployments for autonomous agents have not been defined systematically.
@@
+ It employs STBP (Wu et al., 2018) for learning event-based car data.
@@
Viale et al. (2022)
+ It employs STBP (Wu et al., 2018) for learning event-based road lane data.
@@ 2.2 Motivational Case Study
To highlight the potentials of further optimizing SNNs for autonomous agents from current state-of-the-art, we perform an experimental case study that considers applying different levels of weight quantization (i.e., precision) to an SNN model from the work of Viale et al. (2021) with event-based NCARS dataset. Details of the experimental setup will be discussed further in Section 5. The experimental results shown in Figure 1 show several key observations as discussed in the following.
@@ 2.3 Our Novel Contributions
To address the targeted research problem and the related challenges, we propose a novel framework called SNN4Agents for developing energy-efficient SNNs for autonomous agents. The overview of our SNN4Agents framework is shown in Figure 2 and its key steps are briefly described in the following.
  • Weight Quantization (Section 4.2): To find the appropriate quantization settings that meet the memory constraint, we perform design space exploration for different precision levels while observing their impact on the accuracy.
  • Timestep Reduction (Section 4.3): To find the appropriate processing timesteps that meet the latency constraint, we perform design space exploration for different timestep values while observing their impact on the accuracy.
  • Attention Window Reduction (Section 4.4): To reduce the compute requirements and hence the processing power/energy, we explore different attention window sizes from the input samples and observe their impact on the accuracy.
  • Joint Optimization Strategy (Section 4.5): To maximize the benefits from the previous individual optimization steps, we perform a joint optimization strategy by leveraging design space exploration that considers the appropriate settings from individual optimization steps.
@@ 3.1.1 Overview
Spiking Neural Networks (SNNs) are considered the third generation neural networks Maass (1997). An overview of the SNNs’ functionality is shown in Figure 3. The input spike trains are processed by the spiking neurons and the information is propagated to the neurons in the following layers. Among the most popular spiking neuron models, the Leaky-Integrate-and-Fire (LIF) (Wang et al., 2014) emerges as an efficient trade-off between complexity and plausibility. The operational time of a neuron to process a spike train from a single input data (e.g., an image pixel) is defined as timestep (Putra and Shafique, 2023b). At each timestep, the membrane potential $V_{m}$ of the neuron increases for each incoming input spike. When $V_{m}$ overcomes the voltage threshold $V_{th}$, the neuron emits an output spike Putra et al. (2022b). In this way, the spikes propagate from input to output and multiple layers of spiking neurons in a network.
@@ 3.4 SNN Architecture
Table 2: SNN architecture with 100x100 attention window, based on the work of Viale et al. (2021).
@@
Table 3: SNN architecture with 50x50 attention window, based on the work of Viale et al. (2021).
@@ 4.2 Model Compression through Quantization
To effectively compress the model size, we perform weight quantization through PTQ with TR rounding scheme. To do this, we first train the given network without quantization, while employing baseline settings for timestep, attention window, and training epoch. For this scenario, we employ 32-bit precision, 20 timestep, 100x100 attention window, and 200 training epoch. Once the training phase is finished, we perform quantization process to the trained network. Afterward, we perform DSE under different weight precision levels (i.e., 32, 16, 12, 10, 8, 6, and 4 bit) to evaluate their impact on the accuracy; see the parameter settings for DSE in Table 4.
@@ 4.3 Latency Optimization through Timestep Reduction
To effectively reduce the processing time, we perform timestep reduction. To do this, we shorten the timestep of SNN processing from the baseline setting, thereby curtailing the time window for presenting the spike trains. Here, we consider different timestep settings (i.e., 20, 15, 10, and 5) for exploring their impact on the accuracy. Once we reduce the timestep, the network is trained under baseline setting of precision level (no quantization), attention window, and training epoch. For this scenario, we employ 32-bit precision, 100x100 attention window, and 200 training epoch. If we do not see notable differences from the baseline accuracy, then we may employ a smaller precision level, such as 16- and 4-bit weights, to trigger accuracy variation. The parameter settings for DSE are provided in Table 5.
@@ 4.4 Attention Window Reduction for Input Samples
We also aim at reducing the size of attention window of the input samples to optimize the computational requirements, and hence the latency and energy consumption. To do this, we consider different attention window sizes (i.e., 100x100 and 50x50), while employing baseline settings for precision level, timestep, and training epoch; see the parameter settings for DSE in Table 6.
@@
32 20 100x100 and 50x50 0-200
Experimental results of our DSE are presented in Figure 9. From these results, we observe that accuracy obtained by a smaller attention window (50x50) can saturate faster than a larger one (100x100) thus offering a faster training time; see 6. Meanwhile, a larger attention window offers better accuracy than a smaller one; see 7. The reason is that, a smaller attention window provides less information, hence requiring a shorter time for training the network yet limiting the accuracy that can be achieved. These observations suggest that the reduction of attention window can be employed as long as the targeted accuracy is met.
Figure 9: Results of accuracy across different attention window sizes (i.e., 100x100 and 50x50).
@@ 4.5 Joint Optimization Strategy
  • We perform DSE for the following settings: (1) 10-32 bit precision levels of quantization, (2) 5-20 timesteps, and (3) 50x50 and 100x100 attention window sizes.
@@ 5 Evaluation Methodology
To evaluate our SNN4Agents framework, we build and employ the experimental setup shown in Fig. 10, while considering the same evaluation conditions as widely used in the SNN community for autonomous agents (Putra and Shafique, 2023a; Viale et al., 2021). We use Python-based implementation and run it on Nvidia RTX 6000 Ada GPU machines for evaluating different performance metrics of our SNN4Agents, including accuracy, processing time, and power consumption. Then, the processing time and power consumption are leveraged to estimate the energy consumption. Meanwhile, memory footprint is estimated by leveraging the precision level and the number of weights in the corresponding SNN architecture. For the case of 100x100 attention window, we consider the network architecture from Table 2; while for the case of 50x50, we consider the network architecture from Table 3. For the workload, we consider the NCARS dataset (Sironi et al., 2018). For the DSE purpose, we consider parameter settings shown in Table 7, as well as the state-of-the-art work CarSNN (Viale et al., 2021) as the baseline.
@@
[…]
@@ 6.1 Maintaining High Accuracy
Experimental results for accuracy are presented in Figure 11. These results show that the baseline model (32b_20t) can achieve 85.95% accuracy, while our optimized SNN models with 100x100 attention window achieve 84.12% - 85.76% accuracy and our optimized SNN models with 50x50 attention window achieve 76.81% - 78.56% accuracy. In the 100x100 attention window case, we observe that 10-bit precision levels generally lead to lower learning quality (i.e., accuracy) across all training epochs as compared to other precision levels; see 1 and 2. These indicate that smaller precision levels have smaller value representation capabilities than the larger ones, and hence they require a longer training period to achieve comparable accuracy. For instance, 10-bit precision levels can achieve close to 80% accuracy in 40 training epoch, while other precision levels can achieve it in 20 training epoch. Due to the same reason, similar patterns are observed in the 50x50 attention window case; see 3 and 4. However, in this case, the accuracy scores from different precision levels are not much different, as they lay within around 2% accuracy range; see 4. The reason is that, a smaller attention window means less information to be conveyed and processed by different precision levels, thereby limiting the final accuracy.
Figure 11: Experimental results for accuracy across different parameter settings, including precision levels, timesteps, and training epochs under (a) 100x100 attention window, and (b) 50x50 attention window.
@@ 6.2 Memory Savings
Experimental results for memory footprint are presented in Figure 12. These results show that, in general, our weight quantization step effectively reduces the memory footprint up to 68.75% due to smaller precision levels for representing weight values. For instance, in the 100x100 attention window case, our weight quantization leads to about 50% memory saving for 16b_100w, 62.50% memory saving 12b_100w, and 68.75% memory saving 10b_100w; see 5. Meanwhile, in the 50x50 attention window case, our weight quantization leads to about 91.42% memory saving for 32b_50w, 95.71% memory saving for 16b_50w, 96.78% memory saving 12b_50w, and 97.32% memory saving 10b_50w; see 6. The significant memory savings obtained in the 50x50 attention window case come from the quantization as well as the reduction on the number of weights due to architectural differences in the network as shown in Table 2 and Table 3.
vspace-0.2cm
Figure 12: Experimental results for memory footprint normalized to the baseline (32b_100w) across different precision levels and attention window sizes (i.e., 100x100 and 50x50).
@@ 6.3 Processing Time Speed-Up
Experimental results for processing latency are presented in Figure 13. These results show that, in general, our timestep reduction step effectively reduce the processing latency as compared to the baseline model (32b_20t). For instance, in the 100x100 attention window case, our timestep reduction leads to speed-ups by 1.27x - 1.31x for 15 timestep setting (i.e., 32b_15t, 16b_15t, 12b_15t, and 10b_15t); by 1.86x - 2.02x for 10 timestep setting (i.e., 32b_10t, 16b_10t, 12b_10t, and 10b_10t); as well as by 3.58x - 3.69x for 5 timestep setting (i.e., 32b_5t, 16b_5t, 12b_5t, and 10b_5t); as shown by 7. Meanwhile, in the 50x50 attention window case, our timestep reduction leads to speed-ups by 1.28x - 1.32x for 15 timestep setting (i.e., 32b_15t, 16b_15t, 12b_15t, and 10b_15t); by 2.04x - 2.11x for 10 timestep setting (i.e., 32b_10t, 16b_10t, 12b_10t, and 10b_10t); as well as by 3.85x - 3.95x for 5 timestep setting (i.e., 32b_5t, 16b_5t, 12b_5t, and 10b_5t); as shown by 8. These significant speed-ups come from the reduced timesteps which curtail the processing time for spike trains.
Figure 13: Experimental results for latency normalized to the baseline (32b_20t) across different parameter settings, including precision levels, timesteps, and training epochs under (a) 100x100 attention window, and (b) 50x50 attention window.

### v2

@@ Abstract
Recent trends have shown that autonomous agents, such as Autonomous Ground Vehicles (AGVs), Unmanned Aerial Vehicles (UAVs), and mobile robots, effectively improve human productivity in solving diverse tasks. However, since these agents are typically powered by portable batteries, they require extremely low power/energy consumption to operate in a long lifespan. To solve this challenge, neuromorphic computing has emerged as a promising solution, where bio-inspired Spiking Neural Networks (SNNs) use spikes from event-based cameras or data conversion pre-processing to perform sparse computations efficiently. However, the studies of SNN deployments for autonomous agents are still at an early stage. Hence, the optimization stages for enabling efficient embodied SNN deployments for autonomous agents have not been defined systematically. Toward this, we propose a novel framework called SNN4Agents that consists of a set of optimization techniques for designing energy-efficient embodied SNNs targeting autonomous agent applications. Our SNN4Agents employs weight quantization, timestep reduction, and attention window reduction to jointly improve the energy efficiency, reduce the memory footprint, optimize the processing latency, while maintaining high accuracy. In the evaluation, we investigate use cases of event-based car recognition, and explore the trade-offs among accuracy, latency, memory, and energy consumption. The experimental results show that our proposed framework can maintain high accuracy (i.e., 84.12% accuracy) with 68.75% memory saving, 3.58x speed-up, and 4.03x energy efficiency improvement as compared to the state-of-the-art work for the NCARS dataset. In this manner, our SNN4Agents framework paves the way toward enabling energy-efficient embodied SNN deployments for autonomous agents.
@@ 2 Introduction
In recent years, the interest in implementing neuromorphic artificial intelligence based on Spiking Neural Networks (SNNs) for autonomous agents (so-called SNN-based autonomous agents) has rapidly increased (Putra et al., 2024; Bartolozzi et al., 2022). The reason is that, SNNs can offer high accuracy due to effective learning mechanism (Rathi et al., 2023; Putra and Shafique, 2021b, 2024), low computation latency due to efficient neural/spike coding (Guo et al., 2021), and ultra low power/energy consumption due to sparse spike-based operations (Putra and Shafique, 2020; Schuman et al., 2022). To realize such systems in real life, capabilities of solving machine learning (ML) tasks like image classification (Putra et al., 2021b, 2022a, 2023), object detection (Viale et al., 2021; Cordone et al., 2022), or object segmentation (Li et al., 2022) from images/videos are required. Besides such functionalities, SNN-based autonomous agents also require (1) small memory footprint as they typically employ resource-constrained hardware platforms, (2) low power/energy consumption to preserve the battery lifespan as they are typically powered by portable batteries, and (3) real-time output with high accuracy to provide quick decision (Bonnevie et al., 2021; Putra and Shafique, 2022, 2023b). To maximize the benefits of SNN sparse operations, event-based data can be employed as it directly provides a compatible data format for SNN processing and minimizes the pre-processing stage, such as the data-to-spike conversion (e.g., pixel data to spike train) and the spike coding. Therefore, the developments of SNN-based autonomous agents also need to consider event-based data, such as the NCARS dataset (Sironi et al., 2018; Bano et al., 2024).
@@ 2.1 State-of-the-Art Works and Their Limitations
The study for developing SNN-based autonomous agents is still at an early stage. The state-of-the-art works are summarized in Table 1. Here, we observe that most of the existing works focus on proposing frameworks and/or techniques for achieving high accuracy (Putra and Shafique, 2022, 2023a, 2023b). However, these works have not considered event-based data workloads, therefore requiring a relatively complex pre-processing stage, including data-to-spike conversion and spike coding. Some other works explore techniques for achieving high accuracy considering event-based automotive data (Viale et al., 2021, 2022; Cordone et al., 2022). However, these works have not considered optimizing the model size to fit into the resource-constrained autonomous agents. All the above-discussed limitations of the state-of-the-art expose that, the optimization stages for enabling efficient SNN deployments for autonomous agents have not been defined systematically.
@@
+ It employs the STBP (Wu et al., 2018) for learning event-based car data.
@@
(Viale et al., 2022)
+ It employs the STBP (Wu et al., 2018) for learning event-based road lane data.
@@ 2.2 Motivational Case Study
To highlight the potentials of further optimizing SNNs for autonomous agents from the current state-of-the-art, we perform an experimental case study that considers applying different levels of weight quantization (i.e., precision) to an SNN model from the work of Viale et al. (2021) with event-based NCARS dataset. Details of the experimental setup will be discussed further in Section 5. The experimental results shown in Figure 1 show several key observations as discussed in the following.
@@ 2.3 Our Novel Contributions
To address the targeted research problem and the related challenges, we propose a novel framework called SNN4Agents for developing energy-efficient embodied SNNs for autonomous agents. The overview of our SNN4Agents framework is shown in Figure 2 and its key steps are briefly described in the following.
  • Weight Quantization (Section 4.2): It aims to find the appropriate quantization settings that meet the memory constraint, we perform design space exploration for different precision levels while observing their impact on the accuracy.
  • Timestep Reduction (Section 4.3): It aims to find the appropriate processing timesteps that meet the latency constraint, we perform design space exploration for different timestep values while observing their impact on the accuracy.
  • Attention Window Reduction (Section 4.4): It aims to reduce the compute requirements and hence the processing power/energy, we explore different attention window sizes from the input samples and observe their impact on the accuracy.
  • Joint Optimization Strategy (Section 4.5): It aims to maximize the benefits from the previous individual optimization steps, we perform a joint optimization strategy by leveraging design space exploration that considers the appropriate settings from individual optimization steps.
@@ 3.1.1 Overview
Spiking Neural Networks (SNNs) are considered the third generation neural networks (Maass, 1997). An overview of the SNNs’ functionality is shown in Figure 3. The input spike trains are processed by the spiking neurons and the information is propagated to the neurons in the following layers. Among the most popular spiking neuron models, the Leaky-Integrate-and-Fire (LIF) (Izhikevich, 2004; Wang et al., 2014) emerges as an efficient trade-off between complexity and plausibility. The operational time of a neuron to process a spike train from a single input data (e.g., an image pixel) is defined as timestep (Putra and Shafique, 2023b). At each timestep, the membrane potential $V_{m}$ of the neuron increases for each incoming input spike. When $V_{m}$ overcomes the voltage threshold $V_{th}$, the neuron emits an output spike (Roy et al., 2019; Putra et al., 2021a, 2022b). In this way, the spikes propagate from input to output and multiple layers of spiking neurons in a network.
@@ 3.4 SNN Architecture
Table 2: SNN architecture with 100$\times$100 attention window, based on the work of Viale et al. (2021).
@@
Table 3: SNN architecture with 50$\times$50 attention window, based on the work of Viale et al. (2021).
@@ 4.2 Model Compression through Quantization
To effectively compress the model size, we perform weight quantization through PTQ with TR rounding scheme. To do this, we first train the given network without quantization, while employing baseline settings for timestep, attention window, and training epoch. For this scenario, we employ 32-bit precision, 20 timestep, 100$\times$100 attention window, and 200 training epoch. Once the training phase is finished, we perform quantization process to the trained network. Afterward, we perform DSE under different weight precision levels (i.e., 32, 16, 12, 10, 8, 6, and 4 bit) to evaluate their impact on the accuracy; see the parameter settings for DSE in Table 4.
@@ 4.3 Latency Optimization through Timestep Reduction
To effectively reduce the processing time, we perform timestep reduction. To do this, we shorten the timestep of SNN processing from the baseline settings, thereby curtailing the time window for presenting the spike trains. Here, we consider different timestep settings (i.e., 20, 15, 10, and 5) for exploring their impact on the accuracy. Once we reduce the timestep, the network is trained under baseline settings of precision level (no quantization), attention window, and training epoch. For this scenario, we employ 32-bit precision, 100$\times$100 attention window, and 200 training epoch. If we do not see notable differences from the baseline accuracy, then we may employ a smaller precision level, such as 16- and 4-bit weights, to trigger accuracy variation. The parameter settings for DSE are provided in Table 5.
@@ 4.4 Attention Window Reduction for Input Samples
We also aim at reducing the size of attention window of the input samples to optimize the computational requirements, and hence the latency and energy consumption. To do this, we consider different attention window sizes (i.e., 100$\times$100 and 50$\times$50), while employing baseline settings for precision level, timestep, and training epoch; see the parameter settings for DSE in Table 6.
@@
32 20 100$\times$100 and 50$\times$50 0-200
Experimental results of our DSE are presented in Figure 9. From these results, we observe that accuracy obtained by a smaller attention window (50$\times$50) can saturate faster than a larger one (100x100), thus offering a faster training time; see 6. Meanwhile, a larger attention window offers better accuracy than a smaller one; see 7. The reason is that, a smaller attention window provides less information, thus requiring a shorter time for training the network yet limiting the accuracy that can be achieved. These observations suggest that the reduction of attention window can be employed as long as the targeted accuracy is met.
Figure 9: Results of accuracy across different attention window sizes (i.e., 100$\times$100 and 50$\times$50).
@@ 4.5 Joint Optimization Strategy
  • We perform DSE for the following settings: (1) 10-32 bit precision levels of quantization, (2) 5-20 timesteps, and (3) 50$\times$50 and 100$\times$100 attention window sizes.
@@ 5 Evaluation Methodology
To evaluate our SNN4Agents framework, we build and employ the experimental setup shown in Fig. 10, while considering the same evaluation conditions as widely used in the SNN community for autonomous agents (Putra and Shafique, 2023a; Viale et al., 2021). We use Python-based implementation and run it on Nvidia RTX 6000 Ada GPU machines for evaluating different performance metrics of our SNN4Agents, including accuracy, processing time, and power consumption. Then, the processing time and power consumption are leveraged to estimate the energy consumption. Meanwhile, memory footprint is estimated by leveraging the precision level and the number of weights in the corresponding SNN architecture. For the case of 100$\times$100 attention window, we consider the network architecture from Table 2; while for the case of 50$\times$50, we consider the network architecture from Table 3. For the workload, we consider the NCARS dataset (Sironi et al., 2018). For the DSE purpose, we consider parameter settings shown in Table 7, as well as the state-of-the-art work CarSNN (Viale et al., 2021) as the baseline.
@@
  • $B$b_$W$w: It represents a network with $B$ bit-precision and $W$$\times$$W$ attention window, which encompasses different timestep settings. This term is used when discussing memory footprint in Section 6.2, as networks with different timesteps have the same weight memory footprint.
  • $B$b_$T$t_$W$w: It represents a network with $B$ bit-precision, $T$ timestep, and $W$$\times$$W$ attention window. This term is used when discussing the trade-off analysis in Section 6.5.
  • $T$t_$W$w: It represents a network with $T$ timestep and $W$$\times$$W$ attention window. This term is used when discussing the computational complexity in Section 6.6.
@@ 6.1 Maintaining High Accuracy
Experimental results for accuracy are presented in Figure 11. These results show that the baseline model (32b_20t) can achieve 85.95% accuracy, while our optimized SNN models with 100$\times$100 attention window achieve 84.12% - 85.76% accuracy and our optimized SNN models with 50$\times$50 attention window achieve 76.81% - 78.56% accuracy. In the 100$\times$100 attention window case, we observe that 10-bit precision levels generally lead to lower learning quality (i.e., accuracy) across all training epochs as compared to other precision levels; see 1 and 2. These indicate that smaller precision levels have smaller value representation capabilities than the larger ones, and hence they require a longer training period to achieve comparable accuracy. For instance, 10-bit precision levels can achieve close to 80% accuracy in 40 training epoch, while other precision levels can achieve it in 20 training epoch. Due to the same reason, similar patterns are observed in the 50$\times$50 attention window case; see 3 and 4. However, in this case, the accuracy scores from different precision levels are not much different, as they lay within 2% accuracy range; see 4. The reason is that, a smaller attention window means less information to be conveyed and processed by different precision levels, thereby limiting the final accuracy.
Figure 11: Experimental results for accuracy across different parameter settings, including precision levels, timesteps, and training epochs under (a) 100$\times$100 attention window, and (b) 50$\times$50 attention window.
@@ 6.2 Memory Savings
Experimental results for memory footprint are presented in Figure 12. These results show that, in general, our weight quantization step effectively reduces the memory footprint up to 68.75% due to smaller precision levels for representing weight values. For instance, in the 100$\times$100 attention window case, our weight quantization leads to about 50% memory saving for 16b_100w, 62.50% memory saving 12b_100w, and 68.75% memory saving 10b_100w; see 5. Meanwhile, in the 50$\times$50 attention window case, our weight quantization leads to about 91.42% memory saving for 32b_50w, 95.71% memory saving for 16b_50w, 96.78% memory saving 12b_50w, and 97.32% memory saving 10b_50w; see 6. The significant memory savings obtained in the 50$\times$50 attention window case come from the quantization as well as the reduction on the number of weights due to architectural differences in the network as shown in Table 2 and Table 3.
Figure 12: Experimental results for memory footprint normalized to the baseline (32b_100w) across different precision levels and attention window sizes (i.e., 100$\times$100 and 50$\times$50).
@@ 6.3 Processing Time Speed-Up
Experimental results for processing latency are presented in Figure 13. These results show that, in general, our timestep reduction step effectively reduce the processing latency as compared to the baseline model (32b_20t). For instance, in the 100$\times$100 attention window case, our timestep reduction leads to speed-ups by 1.27x - 1.31x for timestep=15 (i.e., 32b_15t, 16b_15t, 12b_15t, and 10b_15t); by 1.86x - 2.02x for timestep=10 (i.e., 32b_10t, 16b_10t, 12b_10t, and 10b_10t); as well as by 3.58x - 3.69x for timestep=5 (i.e., 32b_5t, 16b_5t, 12b_5t, and 10b_5t); as shown by 7. Meanwhile, in the 50$\times$50 attention window case, our timestep reduction leads to speed-ups by 1.28x - 1.32x for timestep=15 (i.e., 32b_15t, 16b_15t, 12b_15t, and 10b_15t); by 2.04x - 2.11x for timestep=10 (i.e., 32b_10t, 16b_10t, 12b_10t, and 10b_10t); as well as by 3.85x - 3.95x for timestep=5 (i.e., 32b_5t, 16b_5t, 12b_5t, and 10b_5t); as shown by 8. These significant speed-ups come from the reduced timesteps which curtail the processing time for spike trains.
Figure 13: Experimental results for latency normalized to the baseline (32b_20t) across different parameter settings, including precision levels, timesteps, and training epochs under (a) 100$\times$100 attention window, and (b) 50$\times$50 attention window.